OpenStudy (anonymous):

What is the standard form equation of the line passing through (5, -3) with a slope of -3?

OpenStudy (anonymous):

the standard form equation is y=mx+b m is the slope, and b is the y intercept (the y number at which x=0)

OpenStudy (anonymous):

Use point-slope form first.

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

You already have the slope, so the only thing you need to find is the y intercept

OpenStudy (anonymous):

plug in your slope and plug in your point as (x1,y1) then just solve for y. :)

OpenStudy (anonymous):

3x + y = -12?

OpenStudy (anonymous):

solve for y

OpenStudy (anonymous):

i don't need to though, that would be the answer. correct?

OpenStudy (anonymous):

yes. :) I believe so.